# Should I Share this Translation? Evaluating Quality Feedback for User Reliance on Machine Translation

### v2

@@
As people increasingly use AI systems in work and daily life, feedback mechanisms that help them use AI responsibly are urgently needed, particularly in settings where users are not equipped to assess the quality of AI predictions. We study a realistic Machine Translation (MT) scenario where monolingual users decide whether to share an MT output, first without and then with quality feedback. We compare four types of quality feedback: explicit feedback that directly give users an assessment of translation quality using (1) error highlights and (2) LLM explanations, and implicit feedback that helps users compare MT inputs and outputs through (3) backtranslation and (4) question–answer (QA) tables. We find that all feedback types, except error highlights, significantly improve both decision accuracy and appropriate reliance. Notably, implicit feedback, especially QA tables, yields significantly greater gains than explicit feedback in terms of decision accuracy, appropriate reliance, and user perceptions – receiving the highest ratings for helpfulness and trust, and the lowest for mental burden. [^1] https://github.com/dayeonki/mt_quality_feedback
@@ 1 Introduction
Рис. 1: Overview of our study setup. In our human study, each English-speaking monolingual participant reviews a sequence of 20 decision-making examples. Each example is shown in a two-step process: ➊ Independent decision-making: Participants first make judgments based solely on the English source and its Spanish MT output and ➋ AI-Assisted decision-making: They then reassess the same example with one of four randomly assigned feedback types . For each step, they respond to two questions: (1) Shareability: To the best of your knowledge, is the Spanish translation good enough to safely share with your Spanish-speaking neighbor? and (2) Confidence: How confident are you in your assessment?
@@
Artificial Intelligence (AI) are increasingly deployed to support human decision-making across a wide range of domains (Buçinca et al., 2021; Dastin, 2022; Ma et al., 2024). As these systems are adopted by the general public, there is a growing need for feedback mechanisms that help users construct their own functional explanations – reasoning grounded in the goals and consequences of an AI output to determine how and when to rely on it for safe and effective use (Lombrozo and Wilkenfeld, 2019; Schoeffer et al., 2024). Prior work has evaluated various forms of feedback through human studies, such as error highlights (Khashabi et al., 2018; Carton et al., 2020) or free-text explanations (Bussone et al., 2015; Bansal et al., 2021; Buçinca et al., 2021). However, many human-centered studies evaluating the impact of feedback in real-world application settings are still needed. Although designing such studies is challenging, as it requires accounting for the knowledge and assumptions people bring to decision-making tasks (Lage et al., 2019; Wiegreffe and Marasovic, 2021), it remains essential to adopt a human-centered evaluation since these feedback mechanisms are ultimately intended to support human users (Boyd-Graber et al., 2022; Zhu et al., 2024).
@@
To address this, and in line with prior work that empirically investigates real-world use cases in other domains (Hong et al., 2020; Bhatt et al., 2020; Liao et al., 2020), we evaluate AI feedback for Machine Translation (MT), where many users critically need support because they lack the language proficiency needed to evaluate MT outputs. Imagine a situation during the COVID-19 pandemic: You regularly read official guidelines in English, but your Spanish-speaking neighbor cannot access this information. You turn to MT to share this information – but as an English monolingual, how can you determine whether the Spanish MT output is accurate enough to safely share, or if it contains critical errors risking misinformation? This is a practical yet challenging scenario for monolingual users, who lack both source language proficiency (Bowker and Ciro, 2019; Liebling et al., 2020) and domain expertise (Nourani et al., 2020; Lee and Chew, 2023) to reliably evaluate MT quality.
@@
We conduct a human study with 73 English-speaking monolingual participants, where they are asked to decide whether Spanish MT outputs are safe to share with a hypothetical Spanish-speaking neighbor. For each of 20 examples, participants first make a decision without feedback (Independent step), then reassess the same example with a randomly assigned form of feedback (AI-Assisted step), as illustrated in Figure 1.
Our findings reveal that all feedback types except error highlights significantly improve both decision accuracy and appropriate reliance (§4.1). Implicit feedback generally outperforms explicit types: backtranslation yields significantly higher appropriate reliance than error highlights, and QA table leads to significantly higher gains in both metrics than both explicit feedback types (§4.2). While explicit feedback prompts more decision changes, it also results in higher over-reliance, which are cases where participants change from correct to incorrect decisions after viewing feedback (§4.3). We also find that participants are better at recognizing good translations than identifying problematic ones (§5.1). In terms of self-reported perceptions on mental burden, helpfulness, and trust, QA table feedback consistently receives the best ratings (§5.2). Finally, we present participants’ responses to identify which aspects of each quality feedback they found to be helpful (§5.3).
@@ 2.1 Reliance on AI Systems
One of the core challenge in human-AI collaboration is achieving appropriate reliance, which is accepting correct AI advice while rejecting incorrect advice (Eckhardt et al., 2024). In contrast, under-reliance (rejecting correct advice) and over-reliance (accepting incorrect advice) are both undesirable. We build our work on this line of measuring user reliance in AI systems, but more specifically in the context of Machine Translation (MT).
@@ 2.3 Impact of Quality Feedback on Reliance on MT
Various forms of quality feedback have been proposed, including quality estimation (QE) score, backtranslation (Somers, 2005), error highlights that flag problematic spans in the MT output (Eksi et al., 2021; Rubino et al., 2021), textual explanations of metric outputs (Fomicheva et al., 2022; Xu et al., 2023; Jiang et al., 2024; Lu et al., 2024), and question–answer (QA) pairs designed to indicate potential errors in the translation (Sugiyama et al., 2015; Krubiński et al., 2021; Han et al., 2022; Ki et al., 2025; Fernandes et al., 2025).
Several human studies have evaluated quality feedback on user decision-making and reliance on MT. For example, Zouhar et al. (2021) show that backtranslation significantly increased user confidence in translations, even when it did not improve decision accuracy. Mehandru et al. (2023) demonstrate that backtranslation can help users detect critical errors more effectively than QE scores in clinical settings. However, results across these studies remain mixed, and comparisons with more recent feedback mechanisms are still lacking. Our work aims to address this gap.
@@ 3 Methods
Рис. 2: During the AI-assisted decision-making step, each participant is presented with an English source, Spanish translation, and one of four randomly assigned quality feedback types. For error highlights, we also show a color-coded legend ( Minor | Major | Critical) and for QA table, answer texts are displayed in orange when they are identical or highly similar, else, blue.
@@ 3.1 Study Design
Specifically, we situate participants in a scenario where an English monolingual speaker reads official COVID-19 guidelines in English and decide whether the Spanish MT output is of sufficient quality to safely share with a Spanish-speaking neighbor. Each example is presented in two steps: ➊ Independent step, where participants first make judgments based solely on the English source and its Spanish translation, and a subsequent ➋ AI-Assisted step, where the same example is paired with a specific type of quality feedback to support decision-making. Examples are shown in randomized order, and two attention checks were included.
@@
For each example, participants are first asked to make judgments based solely on an English sentence and its corresponding Spanish translation, without any quality feedback. They are asked to answer two questions: (1) Shareability: To the best of your knowledge, is the Spanish translation good enough to safely share with your Spanish-speaking neighbor? (with binary options: [Uncaptioned image] Safe to share as-is, and [Uncaptioned image] Needs bilingual review before sharing); and (2) Confidence: How confident are you in your assessment? (on a five-point Likert scale from 1:Very Unconfident to 5:Very Confident). Since the recruited participants are monolingual English speakers, they are instructed to do their best in assessing the shareability of the MT outputs, despite not being fluent in Spanish.
@@
Subsequently, participants are paired with one of four randomly assigned quality feedback, serving as treatment conditions. They answer the same two questions: shareability and confidence.
@@
Before starting the main study, each participant is asked to answer four pre-task questions regarding their first language, proficiency in English and Spanish, and frequency of using AI translation tools in daily life or work. After completing the main study, each participant answers three post-task questions about their experience with the randomly assigned quality feedback in terms of perceived mental burden, helpfulness, and trust for future use (Hoffman et al., 2019). Detailed descriptions are provided in Appendix C.1.
3.2 Types of Quality Feedback
Figure 2 illustrates an example of each quality feedback. Error highlights and LLM explanation provide explicit quality assessments of MT output, whereas backtranslation and QA table offer implicit assessments that guide participants to compare MT input and output. Details on how each feedback is shown to participants are in Appendix C.1. To control for feedback quality, we balance the error rates of feedback predictions across all types (Appendix D). [^2] We compare the computational and time efficiency for generating each feedback in Appendix E.
@@ 3.3 Stimuli Collection
Рис. 3: Average decision accuracy (left) and CWA (right) for each type of quality feedback. Paired-sample $t$-tests are performed to compare independent and AI-assisted performance and linear mixed-effects ANOVA with Bonferroni corrections to compare different treatment conditions. *: significant with $p$-value < 0.05; **: $p$ < 0.01; ***: $p$ < 0.001; Non-marked ones are not statistically significant. Detailed results are provided in Appendix G.1.
@@ 3.4 Participants
We recruited 73 participants residing in the United States who self-identified English as their first, primary, and fluent language. Recruitment was conducted in two stages to exclude participants proficient in Spanish: (1) A pre-screening survey, where participants reported their English and Spanish proficiency on a five-point scale; and (2) The main task, limited to those who reported high English and low Spanish proficiency in the pre-screening phase. Each main task participant received 5 USD for completing the task (equivalent to 20 USD/hour), [^9] The task took a median of 15 minutes to complete. and 25 participants who achieved over 70% overall decision accuracy received an additional 2 USD bonus. Our institution’s IRB approved to conduct the study. Participants provided informed consent prior to the study. Further details are in Appendix C.2.
Of the 73 participants, 72 reported English as their first language, and one reported both Filipino and English. The average self-reported English proficiency was 5/5 and Spanish proficiency was 1.83/5 . [^10] We detail language proficiency scale in Appendix C. Reported monthly MT usage varied: 3 participants (4.10%) never used MT, 19 (26.0%) rarely used it, 25 (34.2%) used it sometimes, 16 (21.9%) often, and 10 (13.7%) used MT almost every day. Data from one participant who failed both shareability checks was excluded from analysis. Participants were randomly assigned to one of four feedback conditions, with 18 in each group.
@@ 3.5 Dependent Variables
$\mathbf{DecisionAcc.}(E)=\frac{1}{E}\sum_{e\in E}\mathbbm{1}(\hat{s}=s^{*})$ (1)
@@
$\mathbf{CWA}(E)=\frac{1}{E}\sum_{e\in E}\mathrm{sign}(\hat{s})\cdot\frac{c}{5}$ (2)
@@
Switch percentage is a widely used behavioral measure of reliance, capturing how often participants change their decisions after viewing AI feedback (Srivastava et al., 2022; He et al., 2023b). In our context, it reflects how quality feedback influences final shareability judgments (Eckhardt et al., 2024). We compute three metrics following the framework of Schemmer et al. (2023): (1) Over-reliance: the proportion of cases where a participant changes from a correct to an incorrect decision after feedback; (2) Under-reliance: the proportion of cases where a participant does not change from an incorrect decision to a correct one after the quality feedback; (3) Appropriate reliance: the proportion of cases where a participant either corrects an incorrect decision after receiving feedback (switch) or maintains a correct decision (no switch). To account for the differing consequences of reliance depending on shareability, we further break down each reliance metric by shareability label, as detailed in Appendix G.2.
@@ 4 Results
Рис. 4: Average decision accuracy (left) and CWA (right) for each type of quality feedback and shareability label. $\mathbf{n}$ indicates the number of examples aggregated for each condition and label. Independent aggregates responses made without quality feedback across all conditions. **: statistically significant with $p$-value < 0.01; ***: $p$ < 0.001; Non-marked ones are not statistically significant. Detailed results are provided in Appendix G.3.
@@ 4.1 RQ1: Does Quality Feedback Improve MT Decision-Making?
We perform paired-sample $t$-tests to compare independent and AI-assisted performance. As shown in Figure 3, participants generally exhibit higher decision accuracy (left) and appropriate reliance, measured by CWA (right) in the AI-assisted decision-making step compared to the independent step. We observe statistically significant gains in average decision accuracy for LLM explanation (9.31%; p < 0.05), backtranslation (18.7%; p < 0.001), and QA table feedback (16.1%; p < 0.001), resulting in an overall average improvement of 8.32% across all conditions. CWA improves across all conditions with greater extent, averaging 15.3%, indicating that providing any quality feedback is more effective at facilitating participants make accurate decisions with well-calibrated confidence than at improving decision accuracy alone.
In Appendix F, we further show that our two-step setup (➊ Independent → ➋ AI-assisted) does not induce learning effects. The observed gains in decision accuracy and appropriate reliance are attributable to the quality feedback, not repetition.

### v3

@@
As people increasingly use AI systems in work and daily life, mechanisms that help them use AI responsibly are urgently needed, especially when they are not equipped to verify AI predictions themselves. We study a realistic Machine Translation (MT) scenario where monolingual users decide whether to share an MT output, first without and then with quality feedback. We compare four types of quality feedback: explicit feedback that directly give users an assessment of translation quality using (1) error highlights and (2) LLM explanations, and implicit feedback that helps users compare MT inputs and outputs through (3) backtranslation and (4) question–answer (QA) tables. We find that all feedback types, except error highlights, significantly improve both decision accuracy and appropriate reliance. Notably, implicit feedback, especially QA tables, yields significantly greater gains than explicit feedback in terms of decision accuracy, appropriate reliance, and user perceptions – receiving the highest ratings for helpfulness and trust, and the lowest for mental burden. [^1] https://github.com/dayeonki/mt_quality_feedback
@@ 1 Introduction
Рис. 1: Overview of our study setup. In our human study, each English-speaking monolingual participant reviews a sequence of 20 decision-making examples. Each example is shown in a two-step process: ➊ Independent decision-making: Participants first make judgments based solely on the English source and its Spanish MT output and ➋ AI-Assisted decision-making: They then reassess the same example with one of five randomly assigned conditions (one control and four treatments) . For each step, they respond to two questions: (1) Shareability: To the best of your knowledge, is the Spanish translation good enough to safely share with your Spanish-speaking neighbor? and (2) Confidence: How confident are you in your assessment?
@@
Artificial Intelligence (AI) systems are increasingly deployed to support human decision-making across a wide range of domains (Buçinca et al., 2021; Dastin, 2022; Ma et al., 2024). As these systems are adopted by the general public, there is a growing need for feedback mechanisms that help users construct their own functional explanations – reasoning grounded in the goals and consequences of an AI output to determine how and when to rely on it for safe and effective use (Lombrozo and Wilkenfeld, 2019; Schoeffer et al., 2024). Prior work has evaluated various forms of feedback through human studies, such as error highlights (Khashabi et al., 2018; Carton et al., 2020) or free-text explanations (Bussone et al., 2015; Bansal et al., 2021; Buçinca et al., 2021). However, many human-centered studies evaluating the impact of feedback in real-world application settings are still needed. Although designing such studies is challenging, as it requires accounting for the knowledge and assumptions people bring to decision-making tasks (Lage et al., 2019; Wiegreffe and Marasovic, 2021), it remains essential to adopt a human-centered evaluation since these feedback mechanisms are ultimately intended to support human users (Boyd-Graber et al., 2022; Zhu et al., 2024; Carpuat et al., 2025).
@@
To address this, and in line with prior work that empirically investigates real-world use cases in other domains (Hong et al., 2020; Bhatt et al., 2020; Liao et al., 2020), we evaluate AI feedback for Machine Translation (MT), where many users critically need support because they lack the language proficiency needed to evaluate MT outputs. Imagine a situation during the COVID-19 pandemic: You regularly read official guidelines in English, but your Spanish-speaking neighbor cannot access this information. You turn to MT to share this information – but as an English monolingual, how can you determine whether the Spanish MT output is accurate enough to safely share, or if it contains critical errors risking misinformation? This is a practical yet challenging scenario for monolingual users, who lack both source language proficiency (Bowker and Ciro, 2019; Liebling et al., 2020) and domain expertise (Nourani et al., 2020; Lee and Chew, 2023) to reliably evaluate MT quality, and who often lack effective strategies for deciding when to trust imperfect MT (Xiao et al., 2025b).
@@
We conduct a between-subjects human study with 91 English-speaking monolingual participants, where they are asked to decide whether Spanish MT outputs are safe to share with a hypothetical Spanish-speaking neighbor. For each of 20 examples, participants first make a decision (Independent step), then reassess the same example with a randomly assigned condition (AI-Assisted step), as illustrated in Figure 1.
Our findings reveal that all quality feedback interventions except error highlights significantly improve both decision accuracy and appropriate reliance (§4.1). Implicit feedback generally outperforms explicit types: backtranslation yields significantly higher appropriate reliance than error highlights, and QA table leads to significantly higher gains in both metrics than both explicit feedback types (§4.2). While explicit feedback prompts more decision changes, it also results in higher over-reliance, which are cases where participants change from correct to incorrect decisions after viewing feedback (§4.3). We also find that participants are better at recognizing good translations than identifying problematic ones (§5.1). In terms of self-reported perceptions on mental burden, helpfulness, and trust, QA table feedback consistently receives the best ratings (§5.2). Finally, we present participants’ responses to identify which aspects of each quality feedback they found to be helpful (§5.3).
@@ 2.1 Reliance on AI Systems
One of the core challenges in human-AI collaboration is achieving appropriate reliance, which is accepting correct AI advice while rejecting incorrect advice (Eckhardt et al., 2024). In contrast, under-reliance (rejecting correct advice) and over-reliance (accepting incorrect advice) are both undesirable. We build our work on this line of measuring user reliance in AI systems, but more specifically in the context of Machine Translation (MT).
@@ 2.3 Impact of Quality Feedback on Reliance on MT
Various forms of quality feedback have been proposed, including quality estimation (QE) score, backtranslation (Agrawal et al., 2022), error highlights that flag problematic spans in the MT output (Eksi et al., 2021; Rubino et al., 2021; Briakou et al., 2023), textual explanations of metric outputs (Fomicheva et al., 2022; Xu et al., 2023; Jiang et al., 2024; Lu et al., 2024), and question–answer (QA) pairs designed to indicate potential errors in the translation (Sugiyama et al., 2015; Krubiński et al., 2021; Han et al., 2022; Ki et al., 2025; Fernandes et al., 2025).
Only a few human studies have evaluated quality feedback on user decision-making and reliance on MT. For example, Zouhar et al. (2021) show that backtranslation significantly increased user confidence in translations, even when it did not improve decision accuracy. Mehandru et al. (2023) demonstrate that backtranslation can help users detect critical errors more effectively than QE scores in clinical settings. However, results across these studies remain mixed, and comparisons with more recent feedback mechanisms are still lacking. Our work aims to address this gap.
@@ 3 Methods
Рис. 2: During the AI-assisted decision-making step, each treatment group participant is presented with an English source, Spanish translation, and one of four randomly assigned quality feedback types. For error highlights, we also show a color-coded legend ( Minor | Major | Critical) and for QA table, answer texts are displayed in orange when they are identical or highly similar, else, blue.
@@ 3.1 Study Design
Specifically, we situate participants in a scenario where an English monolingual speaker reads official COVID-19 guidelines in English and decide whether the Spanish MT output is of sufficient quality to safely share with a Spanish-speaking neighbor. Each example is presented in two steps: ➊ Independent step, where participants first make judgments based solely on the English source and its Spanish translation, and a subsequent ➋ AI-Assisted step, where the same example is either shown again (control condition) or paired with a specific type of quality feedback to support decision-making (treatment condition). Examples are presented in randomized order, and two attention checks were included.
@@
For each example, participants are first asked to make judgments based solely on an English sentence and its corresponding Spanish translation. They are asked to answer two questions: (1) Shareability: To the best of your knowledge, is the Spanish translation good enough to safely share with your Spanish-speaking neighbor? (with binary options: [Uncaptioned image] Safe to share as-is, and [Uncaptioned image] Needs bilingual review before sharing); and (2) Confidence: How confident are you in your assessment? (on a five-point Likert scale from 1:Very Unconfident to 5:Very Confident). Since the recruited participants are monolingual English speakers, they are instructed to do their best in assessing the shareability of the MT outputs, despite not being fluent in Spanish.
@@
Subsequently, participants are randomly assigned to one of five conditions: a control condition or one of four treatment conditions, each involving a different type of quality feedback (§3.2). Those in the control condition view the same 20 examples twice in succession without receiving any quality feedback. In all conditions, participants answer the same two questions: shareability and confidence.
@@
Before starting the main study, each participant is asked to answer four pre-task questions regarding their first language, proficiency in English and Spanish, and frequency of using AI translation tools in daily life or work. After completing the main study, each participant answers three post-task questions about their experience with the randomly assigned condition in terms of perceived mental burden, helpfulness, and trust for future use (Hoffman et al., 2019). Detailed descriptions are provided in Appendix C.1.
3.2 Types of Quality Feedback Intervention
Figure 2 illustrates an example of each of four quality feedback interventions. Error highlights and LLM explanation provide explicit quality assessments of MT output, whereas backtranslation and QA table offer implicit assessments that guide participants to compare MT input and output. Details on how each feedback is shown to participants are in Appendix C.1. To control for feedback quality, we balance the error rates of feedback predictions across all types (Appendix D). [^2] We compare the computational and time efficiency for generating each feedback in Appendix E.
@@ 3.3 Stimuli Collection
Рис. 3: Average decision accuracy (left) and CWA (right) for each condition. Paired-sample $t$-tests are performed to compare independent and AI-assisted performance and linear mixed-effects ANOVA with Bonferroni corrections to compare different treatment conditions. *: significant with $p$-value < 0.05; **: $p$ < 0.01; ***: $p$ < 0.001; Non-marked: not statistically significant. Detailed results are provided in Appendix F.1.
@@ 3.4 Participants
We recruited 91 participants residing in the United States who self-identified English as their first, primary, and fluent language. Recruitment was conducted in two stages to exclude participants proficient in Spanish: (1) A pre-screening survey, where participants reported their English and Spanish proficiency on a five-point scale; and (2) The main task, limited to those who reported high English and low Spanish proficiency in the pre-screening phase. Each main task participant received 5 USD for completing the task (equivalent to 20 USD/hour), [^9] The task took a median of 14 minutes to complete. and 30 participants who achieved over 70% overall decision accuracy received an additional 2 USD bonus. Our institution’s IRB approved to conduct the study. Participants provided informed consent prior to the study. Further details are in Appendix C.2.
Of the 91 participants, 90 reported English as their first language, and one reported both Filipino and English. The average self-reported English proficiency was 5/5 and Spanish proficiency was 1.83/5.101010We detail language proficiency scale in Appendix C. Reported monthly MT usage varied: 5 participants (5.49%) never used MT, 24 (26.4%) rarely used it, 32 (35.2%) used it sometimes, 19 (20.9%) often, and 11 (12.1%) used MT almost every day. Data from one participant who failed both shareability checks was excluded from analysis. Participants were randomly assigned to one of five conditions, with 18 in each group.
Рис. 4: Average decision accuracy (left) and CWA (right) for each type of quality feedback and shareability label. $\mathbf{n}$ indicates the number of examples aggregated for each condition and label. Independent aggregates responses made without quality feedback across all conditions. **: statistically significant with $p$-value < 0.01; ***: $p$ < 0.001; Non-marked ones are not statistically significant. Detailed results are provided in Appendix F.3.
@@ 3.5 Dependent Variables
$\mathbf{DecisionAcc.}(E)=\frac{1}{|E|}\sum_{e\in E}\mathbbm{1}(\hat{s}=s^{*})$ (1)
@@
$\mathbf{CWA}(E)=\frac{1}{|E|}\sum_{e\in E}\mathrm{sign}(\hat{s})\cdot\frac{c}{5}$ (2)
@@
Switch percentage is a widely used behavioral measure of reliance, capturing how often participants change their decisions after viewing AI feedback (Srivastava et al., 2022; He et al., 2023b). In our context, it reflects how quality feedback influences final shareability judgments (Eckhardt et al., 2024). We compute three metrics following the framework of Schemmer et al. (2023): (1) Over-reliance: the proportion of cases where a participant changes from a correct to an incorrect decision after feedback; (2) Under-reliance: the proportion of cases where a participant does not change from an incorrect decision to a correct one after the quality feedback; (3) Appropriate reliance: the proportion of cases where a participant either corrects an incorrect decision after receiving feedback (switch) or maintains a correct decision (no switch). To account for the differing consequences of reliance depending on shareability, we further break down each reliance metric by shareability label, as detailed in Appendix F.2.
@@ 4.1 RQ1: Does Quality Feedback Improve MT Decision-Making?
We perform paired-sample $t$-tests to compare independent and AI-assisted performance. As shown in Figure 3, participants in all four treatment conditions generally exhibit higher decision accuracy (left) and appropriate reliance, measured by CWA (right) in the AI-assisted decision-making step compared to the independent step. We observe statistically significant gains in average decision accuracy for LLM explanation (9.31%; p < 0.05), backtranslation (18.7%; p < 0.001), and QA table feedback (16.1%; p < 0.001), resulting in an overall average improvement of 8.32% across all conditions. CWA improves across all conditions with greater extent, averaging 15.3%, indicating that providing any quality feedback is more effective at helping participants make accurate decisions with well-calibrated confidence than at improving decision accuracy alone.
We further find that both decision accuracy (M = 0.605, S.E. = 0.036) and CWA (M = 0.172, S.E. = 0.054) in the AI-assisted step are significantly lower in the control condition than in all treatment conditions except for the error highlights group. Moreover, the within-group difference between the independent and AI-assisted steps in the control condition is not statistically significant. This indicates that our two-step setup does not induce learning effects but the observed gains in decision accuracy and appropriate reliance stem from the quality feedback, not from repetition.